OpenStudy (anonymous):

Find the variation constant an equation of variation where y varies inversely as x and y =5 when x=15

OpenStudy (saifoo.khan):

y = k/x k = 15*5 k = 75

OpenStudy (saifoo.khan):

y = 75/x